ate the earned value for the project. All amounts are in Ise the data to caleulate e housands of dollars, ( uration weeka 10 15 ptual Design anaplos וי roductionn otal 10 10 18 umulative B) $100,000 C) $43,250 100 A) $38,850 D) $54,750
Answer #1

Earned Value of the project can be computed by the following formula

%age completion of the project X the cost of completion

In the given case the Earned value can be computed as folllows

Activity %completed Cost to complete $

Earned value $

% completed X cost to complete

Conceptual Design 100% (10,000+6000)=16,000 16,000
Graphics Design 75% (12+6+4)*1000=22,000 16,500
Copyright 5% (4+5+6)*1000=15,000 750
Samples 25% (4+2)*1000=6,000 1500
Tooling 10% (6+9+4)*1000=19,000 1900
Production 10% (2+6+14)*1000=22,000 2200
Total 38850

So, the answer is A)38,850